OpenStudy (anonymous):

MEDAL! write the equation in slope intercept form. What are the slope and y-intercept? -12x+11y=-8

OpenStudy (anonymous):

11y=12x-8 Divide each term in the equation by 11. 11y/11=12x/11-8/11 Simplify y=12x/11-8/11 y=mx+b m=12/11 b=-8/11 m is the slope and b is the y intercept